you have a lump sum of 80000 to invest. you decided to invest in unit trusts.the unit trust aims to perfom at return of cpi which is about 8%,how long will you have to keep these unit trusts if you want your lump sum to double.
Answer by Theo(13342)

f = p * (1+r)^n

f = future value = 160000
p = present value = 80000
note that 160000 = 2 times 80000
r = 8%/100 = .08
n = what you want to find.

your equation is:

160000 = 80000 * (1.08)^n

divide both sides of othis equation by 80000 to get:

2 = 1.08^n

take the natural log of both sides of the equation to get:

ln(2) = ln(1.08^n)

this is equivalent to:

ln(2) = n*ln(1.08)

divide both sides of the equation by ln(1.08) to get:

ln(2) / ln(1.08) = n

solve for n to get:

n = 9.006468.

confirm by replacing n in the original equation with 9.006468.

160000 = 80000 * (1.08)^9.006468

solve to get:

160000 = 160000

this confirms the solution is correct.